OpenStudy (anonymous):

the inverse sine of sin(2pi/3) = 120degrees, correct?

OpenStudy (anonymous):

hell no

OpenStudy (anonymous):

trick question!

OpenStudy (anonymous):

it is a trick to see if you are paying attention

OpenStudy (anonymous):

LOL - whatttttt??? sin-1(sinx)=x!!!

OpenStudy (anonymous):

OH ITS NOT IN THE DAMN RANGE

OpenStudy (anonymous):

ooops. range of arcsine is \[[-\frac{\pi}{2},\frac{\pi}{2}]\]

OpenStudy (anonymous):

BALLS I SAY

OpenStudy (anonymous):

it should be called trickonometry

OpenStudy (anonymous):

yes. 2pi/3*180/pi 2/3*180 2*60 120

OpenStudy (anonymous):

yah you have to find another value whose sine is also \[\frac{2\pi}{3}\]

OpenStudy (anonymous):

so wait - am i right or wrong with my original answer of 120?

OpenStudy (anonymous):

wrongo

OpenStudy (anonymous):

it is \[\frac{\pi}{3}\]

OpenStudy (anonymous):

oh okay - sorry. so how do i find a ....hm. range of arccosine then.....and....

OpenStudy (anonymous):

dang it! i should have guessed...

OpenStudy (anonymous):

\[\sin(\frac{2\pi}{3})=\frac{\sqrt{3}}{2}\] but \[\sin^{-1}(\frac{\sqrt{3}}{2})=\frac{\pi}{3}\]

OpenStudy (anonymous):

so is 60 degrees

OpenStudy (anonymous):

ok think like this for arcsine look on right half of unit circle from \[-\frac{\pi}{2}\] to \[\frac{\pi}{2}\]

OpenStudy (anonymous):

f@ck degrees.

OpenStudy (anonymous):

you are not working in degrees you are not converting degrees to radians forget about degrees

OpenStudy (anonymous):

r the equation machines not working? i can't get mine up

OpenStudy (anonymous):

viagra

OpenStudy (anonymous):

i can't give you another medal man, but i....wood? lol

OpenStudy (anonymous):

for arccosine look at upper half of unit circle. from 2 to pi

OpenStudy (anonymous):

np

OpenStudy (anonymous):

right, i just don't get how i can arbitrarily change the range i'm using for sine...

OpenStudy (anonymous):

wood. right. i am a little slow

OpenStudy (anonymous):

there is nothing arbitrary about it, it is fixed. ok i guess it is arbitrary but it does not change

OpenStudy (anonymous):

and don't get the range of sine confused with the range of arcsine.

OpenStudy (anonymous):

how can i apply the range of the arccosine to an arcsine?

OpenStudy (anonymous):

range of arcsine is \[[-\frac{\pi}{2},\frac{\pi}{2}]\]

OpenStudy (anonymous):

i am sure you know what you mean by that question but i am afraid i do not

OpenStudy (anonymous):

awwww hell. i got it. youdaman

OpenStudy (anonymous):

the range is the range is the range. if you want arccosine of something make sure your answer is between 0 and pi and if you want the arcsine of something your answer must be between -pi/2 andpi/2

OpenStudy (anonymous):

trickonometry. i will remember that one